As a solution of copper sulfate slowly evaporates, beautiful blue crystals made of Cu(II) and sulfate ions form such that water molecules are trapped inside the crystals. The overall formula of the compound is CuSO4 • 5H2O.

i) What is the percent water in this compound

ii) At high temperatures the water in the compound is driven off as steam. What mass percentage of the original sample of the blue solid is lost as a result?
1
Expert's answer
2020-10-12T13:32:58-0400

1)M(Cu) -  63,5 g/mol

M(S) - 32 g/mol

M(O) - 16 g/mol

M(H) - 1 g/ mol

M(CuSO4) =  63.5 + 32 +16*4 = 159.5

M(H2O) = 2 + 16 = 18

M(CuSO4*5H2O) = 159.5 +18 * 5= 249.5

C(%)(H2O) = 18*5/249.5 *100% = 36%

2)Vaporize only water, so 36%